OpenStudy (anonymous):

Simplify the expression. 8n+3-6n/2 A.14n+1/2 B.-n+2 C.n+3/2 D. 5n-3

OpenStudy (anonymous):

Produce n s together and numbers or do the * or / look : -6n/2=-3n Ok ? -6/2=-3 so can produce n s together : -3n+8n=(8-3)=5n we have 3 to but cant produce it so just write that ! Answer is :5n+3